OpenStudy (anonymous):

Evaluate the following expression: (7 + 5b) + 1 , if b = 3. (1 point) 27 23 0 24

OpenStudy (anonymous):

7+5*3+1=7+15+1=23

OpenStudy (anonymous):

yes of course

OpenStudy (anonymous):

20-14=6 6^2=36 36/12=3

OpenStudy (anonymous):

well. ask that, please

OpenStudy (anonymous):

answer is 12h ;180$